# Memorization Sinks: Isolating Memorization during LLM Training

### v1

@@ Appendix E Implementation of Gradient Masking
We generally follow the implementation outlined in (Cloud et al., 2024). We partition each MLP layer into memorization and generalization neurons. We tune this delineation of memorization and generalization neurons by the proportion of generalization neurons $g$. We additionally partition our dataset into examples seen once and the repeated examples. During training, we mask the gradients in each MLP layer such that the gradients from the repeated examples update only a the memorization block, whereas gradients of all other examples are routed to the generalization block.
@@
Min Learning Rate Max Learning $\frac{\text{Max Learning Rate}}{10}$
@@ Appendix F Analysis of Gradient Masking
Like the previous analysis, we will consider a dataset of $N$ $(s_{i},y_{i})$ pairs which are only seen once, and a special $(s_{\textrm{mem}},y_{\textrm{mem}})$ that is repeated in training and we wish to isolate. Thus, the total training set $D_{\textrm{pre}}=\{(s_{i},y_{i})\}_{i=1}^{N}\cup(s_{\textrm{mem}},y_{\textrm{mem}})$.
@@ Appendix G Implementation of MemSinks
Min Learning Rate Max Learning $\frac{\text{Max Learning Rate}}{10}$
@@
We used the same model architecture as reported in Appendix B. We set the first $g$ fraction of neurons in each MLP as the “shared neurons” and left the remaining $1-g$ fraction as the memorization neuron pool. We applied the dropout layer after the GeLU activation function, prior to the downprojection layer.
@@ H.1 Formalization of Training Process
We will treat our data as (embedding, next token) pairs. We consider we have a repeated sequence $\mathbf{s}^{\text{mem}}$ with corresponding next token $\mathbf{e}^{\text{mem}}$. Next, we will assume we have a large dataset of sequences seen only once during training $\mathcal{D}_{\text{once}}=\{(\mathbf{s}^{\text{(1)}},\mathbf{e}^{(1)}),...,(\mathbf{s}^{\text{(N)}},\mathbf{e}^{(N)}\})\}$. For simplicity, we will consider the case where $\forall i$ $\mathbf{e}^{(i)}\neq\mathbf{e}^{\text{mem}}$. Since we treat $\mathbf{W_{\text{proj}}}$ as frozen, we will also define $\epsilon_{\text{shared}}=\min\mathbf{z}(\mathbf{s}^{\text{mem}})_{\text{shared}}^{\top}\mathbf{z}(\mathbf{s}^{(i)})_{\text{shared}}$ and likewise that $\epsilon_{\text{mem}}=\min_{i}\mathbf{z}(\mathbf{s}^{\text{mem}})_{\text{mem}}^{\top}\mathbf{z}(\mathbf{s}^{(i)})_{\text{mem}}$. Intuitively, these quantities lower bound how similar the activations in the shared and memorization neurons are between the repeated example and any other example. For simplicity we will assume that the $||\mathbf{z^{(i)}}||_{2}=1$ for all $\mathbf{z^{(i)}}$ and that the parameter $||\mathbf{W}_{\text{proj}}||_{2}<\frac{C_{\text{proj}}}{2}$ remains bounded throughout training. Finally we assume that the ouput embeddings $e$ are mutually orthogonal.
@@ H.2 Forgetting Under Normal Training Dynamics
Given our assumption that $||\mathbf{z}^{(i)}||_{2}=1$ and the bounded parameter norm assumption $||\mathbf{W}_{\text{proj}}||_{2}<\frac{C_{\text{proj}}}{2}$, it follows that $||\mathbf{W_{\text{proj}}}\mathbf{z}^{(i)}||_{\infty}\leq\frac{C_{V}}{2}$. By Theorem H.1, we have that the entries of $\frac{\exp(-C_{\text{proj}})}{d_{\text{emb}}}\leq\sigma(f(\mathbf{z}^{(i)})\leq\frac{\exp(C_{\text{proj}})}{d_{\text{emb}}-1}$, element wise. In the remainder of the theory, we denote $c_{min}=\frac{\exp(-C_{\text{proj}})}{d_{\text{emb}}}$ and $c_{max}=\frac{\exp(C_{\text{proj}})}{d_{\text{emb}-1}}$.
@@
Next, we will show that the seqTD accumulates memorization in the memorization neurons, as formalized in the following theorem. This theorem also crystalizes some key quantities relating to gradient interference. First of all, we see that the forgetting depends on the number of further gradient steps taken after seeing $\mathbf{s}^{\text{mem}}$. Secondly, we observe that the impact of forgetting dynamics is influcned by how similar the activation of neurons are amongst different examples: controlled by $\epsilon$. The first observation immediately suggests that if some neurons were activated less often, then those neurons would be effectively “store” more memorization.
@@ H.3 Analysis of MemSinks
We will first examine $(\mathbf{e^{\text{mem}}})^{\top}f^{(n)}_{gen-only}(\mathbf{z}^{\text{mem}})$. At any point in training, recall that we can upper and lower bound the value $c_{min}\leq(\mathbf{e}^{\text{(i)}})^{\top}\sigma(f(\mathbf{z}^{\text{mem}}))\leq c_{max}$. As such, observe that $(\mathbf{e}^{\text{(i)}})^{\top}\sigma(f(\mathbf{z}^{\text{mem}}))$ received $k$ updates upper bounded by $\gamma(1-c_{min})$ (from the $k$ obervations of $\mathbf{z}^{\text{mem}}$ and $(N-k)$ updates upper bounded by $\gamma\epsilon_{\text{shared}}c_{min}$ (from the remaining $(N-k)$ observations of the $\mathbf{z}^{(i)}$. This yields the desired claim for (1).
Now, for claim (2) observe that the component $(\mathbf{e^{\text{mem}}})^{\top}f^{(n)}_{mem-only}(\mathbf{z}^{\text{mem}})$ receives $k$ updates lower bounded by $(1-c_{max})$ (again, from the $k$ observations of $\mathbf{z}^{\text{mem}}$, but only $p(N-k)$ updates from other observations, which can likewise be lower bounded by $\gamma\epsilon_{\text{mem}}c_{max}$ This immediately implies the desired claim in (2) ∎
This theorem formalizes the notion that memorization “accumulates” in the memorization neurons when they are shielded from the interference of other sequences sufficiently. In our theory, the extent to which this occurs is dependent on two quantities (1) the fraction of non-repeated sequences for which the memorization neurons are active and (2) the similarity of activations of the repeated example and non-repeated example in the memorization neurons. Relative to algorithm design, however, we will generally only have control over $\rho$ and so we will consider $\epsilon_{\text{shared}}=\epsilon_{\text{mem}}$ out of convenience. Our analysis demonstrates that when $\rho$ is set appropriately low. Some calculation demonstrates that when $\rho<\frac{c_{min}}{c_{max}}-\frac{k}{(N-k)}(c_{max}-c_{min})$, then we will have a seperation in the logits of $\mathbf{s}^{\text{mem}}$ where the memorization neurons primarily contain the memorized example.
@@ Appendix I Large Scale Implementation Details
We consider pretraining datasets of 1 and 2 billion parameters (respectively for the 360M model and 1.7B model). The majority of pretraining tokens are sourced from a subset of the SlimPajama dataset (Shen et al., 2024). However, we consider a setting in which we have an under-sampled domain of 5000 TinyStories examples. We ensure that these 5000 TinyStories (Eldan & Li, 2023) documents are included in the pretraining data before filling the remaining tokens from SlimPajamas.
@@
Min Learning Rate Max Learning $\frac{\text{Max Learning Rate}}{10}$
@@
Min Learning Rate Max Learning $\frac{\text{Max Learning Rate}}{10}$
@@ I.1 Implementation of MemSinks
We will release a full implementation of MemSinks at http://github.com/grghosal/MemSinks. Our implementation builds off of the LitGPT library (AI, 2023).
@@
We implemented MemSinks at every hidden MLP layer by partitioning the neurons into generalizing and memorization sink groups. We then apply the mask over the memorization neurons. For efficiency, we used the same masking across all MLP layers. As the SmolLM (Allal et al., 2025) family models uses a GatedMLP implementation, we apply MemSinks in the output-space after the gating.

### v2

@@ Appendix E Implementation of Gradient Masking
We generally follow the implementation outlined in Cloud et al. (2024). We partition each MLP layer into memorization and generalization neurons. We tune this delineation of memorization and generalization neurons by the proportion of generalization neurons $g$. We additionally partition our dataset into examples seen once and the repeated examples. During training, we mask the gradients in each MLP layer such that the gradients from the repeated examples update only a the memorization block, whereas gradients of all other examples are routed to the generalization block.
@@
Min Learning Rate $\frac{\text{Max Learning Rate}}{10}$
@@ Appendix F Analysis of Gradient Masking
Like the previous analysis, we will consider a dataset of NN $(s_{i},y_{i})$ pairs which are only seen once, and a special $(s_{\textrm{mem}},y_{\textrm{mem}})$ that is repeated in training and we wish to isolate. Thus, the total training set $D_{\textrm{pre}}=\{(s_{i},y_{i})\}_{i=1}^{N}\cup(s_{\textrm{mem}},y_{\textrm{mem}})$.
@@ Appendix G Implementation of MemSinks
Min Learning Rate $\frac{\text{Max Learning Rate}}{10}$
@@
We used the same model architecture as reported in Appendix B. We set the first $g$ fraction of neurons in each MLP as the “shared neurons" and left the remaining 1g1-g fraction as the memorization neuron pool. We applied the dropout layer after the GeLU activation function, prior to the downprojection layer.
@@ H.1 Formalization of Training Process
We will treat our data as (embedding, next token) pairs. We consider we have a repeated sequence $\mathbf{s}^{\text{mem}}$ with corresponding next token $\mathbf{e}^{\text{mem}}$. Next, we will assume we have a large dataset of sequences seen only once during training $\mathcal{D}_{\text{once}}=\{(\mathbf{s}^{\text{(1)}},\mathbf{e}^{(1)}),...,(\mathbf{s}^{\text{(N)}},\mathbf{e}^{(N)}\})\}$. For simplicity, we will consider the case where $\forall i$ $\mathbf{e}^{(i)}\neq\mathbf{e}^{\text{mem}}$. Since we treat $\mathbf{W_{\text{proj}}}$ as frozen, we will also define $\epsilon_{\text{shared}}=\min\mathbf{z}(\mathbf{s}^{\text{mem}})_{\text{shared}}^{\top}\mathbf{z}(\mathbf{s}^{(i)})_{\text{shared}}$ and likewise that $\epsilon_{\text{mem}}=\min_{i}\mathbf{z}(\mathbf{s}^{\text{mem}})_{\text{mem}}^{\top}\mathbf{z}(\mathbf{s}^{(i)})_{\text{mem}}$. Intuitively, these quantities lower bound how similar the activations in the shared and memorization neurons are between the repeated example and any other example. For simplicity we will assume that the $\mathbf{z^{(i)}}||_{2}=1$ for all $\mathbf{z^{(i)}}$ and that the parameter $\mathbf{W}_{\text{proj}}||_{2}<\frac{C_{\text{proj}}}{2}$ remains bounded throughout training. Finally we assume that the ouput embeddings $e$ are mutually orthogonal.
@@ H.2 Forgetting Under Normal Training Dynamics
Given our assumption that $\mathbf{z}^{(i)}||_{2}=1$ and the bounded parameter norm assumption $\mathbf{W}_{\text{proj}}||_{2}<\frac{C_{\text{proj}}}{2}$, it follows that $\mathbf{W_{\text{proj}}}\mathbf{z}^{(i)}||_{\infty}\leq\frac{C_{V}}{2}$. By Theorem H.1, we have that the entries of $\frac{\exp(-C_{\text{proj}})}{d_{\text{emb}}}\leq\sigma(f(\mathbf{z}^{(i)})\leq\frac{\exp(C_{\text{proj}})}{d_{\text{emb}}-1}$, element wise. In the remainder of the theory, we denote $c_{min}=\frac{\exp(-C_{\text{proj}})}{d_{\text{emb}}}$ and $c_{max}=\frac{\exp(C_{\text{proj}})}{d_{\text{emb}-1}}$.
@@
Next, we will show that the seqTD accumulates memorization in the memorization neurons, as formalized in the following theorem. This theorem also crystalizes some key quantities relating to gradient interference. First of all, we see that the forgetting depends on the number of further gradient steps taken after seeing $\mathbf{s}^{\text{mem}}$. Secondly, we observe that the impact of forgetting dynamics is influcned by how similar the activation of neurons are amongst different examples: controlled by $\epsilon$. The first observation immediately suggests that if some neurons were activated less often, then those neurons would be effectively “store" more memorization.
@@ H.3 Analysis of MemSinks
We will first examine $(\mathbf{e^{\text{mem}}})^{\top}f^{(n)}_{gen-only}(\mathbf{z}^{\text{mem}})$. At any point in training, recall that we can upper and lower bound the value $c_{min}\leq(\mathbf{e}^{\text{(i)}})^{\top}\sigma(f(\mathbf{z}^{\text{mem}}))\leq c_{max}$. As such, observe that $(\mathbf{e}^{\text{(i)}})^{\top}\sigma(f(\mathbf{z}^{\text{mem}}))$ received $k$ updates upper bounded by $\gamma(1-c_{min})$ (from the $k$ obervations of $\mathbf{z}^{\text{mem}}$ and (Nk)(N-k) updates upper bounded by $\gamma\epsilon_{\text{shared}}c_{min}$ (from the remaining (Nk)(N-k) observations of the $\mathbf{z}^{(i)}$. This yields the desired claim for (1).
Now, for claim (2) observe that the component $(\mathbf{e^{\text{mem}}})^{\top}f^{(n)}_{mem-only}(\mathbf{z}^{\text{mem}})$ receives $k$ updates lower bounded by $(1-c_{max})$ (again, from the $k$ observations of $\mathbf{z}^{\text{mem}}$, but only p(Nk)p(N-k) updates from other observations, which can likewise be lower bounded by $\gamma\epsilon_{\text{mem}}c_{max}$ This immediately implies the desired claim in (2) ∎
This theorem formalizes the notion that memorization “accumulates" in the memorization neurons when they are shielded from the interference of other sequences sufficiently. In our theory, the extent to which this occurs is dependent on two quantities (1) the fraction of non-repeated sequences for which the memorization neurons are active and (2) the similarity of activations of the repeated example and non-repeated example in the memorization neurons. Relative to algorithm design, however, we will generally only have control over $\rho$ and so we will consider $\epsilon_{\text{shared}}=\epsilon_{\text{mem}}$ out of convenience. Our analysis demonstrates that when $\rho$ is set appropriately low. Some calculation demonstrates that when $\rho<\frac{c_{min}}{c_{max}}-\frac{k}{(N-k)}(c_{max}-c_{min})$, then we will have a seperation in the logits of $\mathbf{s}^{\text{mem}}$ where the memorization neurons primarily contain the memorized example.
@@ Appendix I Large Scale Implementation Details
We consider pretraining datasets of 1 and 2 billion parameters (respectively for the 360M model and 1.7B model). The majority of pretraining tokens are sourced from a subset of the SlimPajama dataset Shen et al. (2024). However, we consider a setting in which we have an under-sampled domain of 5000 TinyStories examples. We ensure that these 5000 TinyStories Eldan & Li (2023) documents are included in the pretraining data before filling the remaining tokens from SlimPajamas.
@@
Min Learning Rate $\frac{\text{Max Learning Rate}}{10}$
@@
Min Learning Rate $\frac{\text{Max Learning Rate}}{10}$
@@ I.1 Implementation of MemSinks
We will release a full implementation of MemSinks at http://github.com/grghosal/MemSinks. Our implementation builds off of the LitGPT library AI (2023).
@@
We implemented MemSinks at every hidden MLP layer by partitioning the neurons into generalizing and memorization sink groups. We then apply the mask over the memorization neurons. For efficiency, we used the same masking across all MLP layers. As the SmolLM Allal et al. (2025) family models uses a GatedMLP implementation, we apply MemSinks in the output-space after the gating.